I am confused!! the answer key in the book is wrong
my problem is:
[9-3/5x]+6=12 this is for absolute values

Answer by sofiyac(983)

|9-3/5x|+6=12 subtract 6 from each side of the equal sign
|9-3/5x|=6
case1
9-3/5x=6 subtract 9 from each side
-3/5x=-3 multiply each side by -5/3
x=-3(-5/3) simplify, 3's cancel out
x=5
case 2
-(9-3/5x)=6
-9+3/5x=6 add 9 to each side
3/5x=15 multiply each side by 5/3
x=15(5/3) simplify
x=25
so possible solutions are x=5 and x=25